OpenStudy (anonymous):

Suppose houses are available for purchase during construction of a new housing development. Two different types of houses are being built: rancher and colonial. The mean price of the ranchers is $319,000 with a standard deviation of $12,000. The mean price of the colonials is $405,000 with a standard deviation of $18,500.

OpenStudy (anonymous):

a) The purchase price for the colonial houses is reduced by $10,000. What would be the new mean for the colonial houses? b) If the purchase price for the colonial houses is reduced by $10,000, how much would the standard deviation for the colonial houses change?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Ahh that makes it easier lol...well lets see.... a) take the new price...add the price of the ranches...divide by 2... so 319,000 + 395,000 ------------------ = ? 2

OpenStudy (johnweldon1993):

And then for b) you need to find the variance...and then take the square root of it to find the new standard deviation... Variance = (Price of ranches)^2 + (new price of colonials)^2 The the square root of that = standard deviation Standard Deviation = \[\large \sqrt{319000^2 + 395000^2}\]

OpenStudy (anonymous):

thank you :)

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (anonymous):

IS B 16,055.72 OR 5,996.09